Skip to main contentSkip to solution

Match List-I with List-II

Let AA and BB be two events such that P(A)=0.2P(A) = 0.2, P(B)=0.4P(B) = 0.4, P(B∣A)=0.5P(B|A) = 0.5

List-IList-II
(A) P(A∩B)P(A \cap B)(I) 0.50.5
(B) P(A∣B)P(A\vert B)(II) 0.80.8
(C) P(A∪B)P(A \cup B)(III) 0.250.25
(D) P(A′)P(A')(IV) 0.10.1

Choose the correct answer from the options given below:

Solution

✅ Correct Option: 1

Given P(A)=0.2P(A) = 0.2, P(B)=0.4P(B) = 0.4, P(B∣A)=0.5P(B|A) = 0.5

Using the conditional probability formula:

P(B∣A)=P(A∩B)P(A)P(B|A) = \dfrac{P(A \cap B)}{P(A)}

P(A∩B)=P(B∣A)×P(A)P(A \cap B) = P(B|A) \times P(A)

=0.5×0.2= 0.5 \times 0.2

=0.1= 0.1

⇒\Rightarrow Matches (IV)


P(A∣B)=P(A∩B)P(B)P(A|B) = \dfrac{P(A \cap B)}{P(B)}

=0.10.4= \dfrac{0.1}{0.4}

=0.25= 0.25

⇒\Rightarrow Matches (III)


P(A∪B)=P(A)+P(B)−P(A∩B)P(A \cup B) = P(A) + P(B) - P(A \cap B)

=0.2+0.4−0.1= 0.2 + 0.4 - 0.1

=0.5= 0.5

⇒\Rightarrow Matches (I)


P(A′)=1−P(A)P(A') = 1 - P(A)

=1−0.2= 1 - 0.2

=0.8= 0.8

⇒\Rightarrow Matches (II)


List-IValueList-II
(A) P(A∩B)P(A \cap B)0.10.1(IV)
(B) $P(AB)$0.250.25
(C) P(A∪B)P(A \cup B)0.50.5(I)
(D) P(A′)P(A')0.80.8(II)

Correct Answer: (A)-(IV), (B)-(III), (C)-(I), (D)-(II)

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question